I need to show steps in solving for the following problem. I cannot figure out how to type (sub 1) after a variable, so I will just write it after the variable in the equation.
y-y sub1 = m(x-x sub 1) ; solve for x

Answer by Alan3354(69443)

y-y sub1 = m(x-x sub 1) ; solve for x
---------
y-y1 = m(x-x1)
x-x1 = (y-y1)/m
x = x1+ (y-y1)/m